# Schema Generation for Large Knowledge Graphs Using Large Language Models

### v1

@@ 7 Conclusion
The size of the dataset is constrained by limited resources available for its construction. Creating high-quality shape schemas is a time-consuming process that requires multiple rounds of refinement by volunteer knowledge engineers. As a result, the current size of the dataset makes it difficult to convincingly evaluate the effectiveness or efficiency of traditional machine learning or deep learning approaches—such as embedding-based models—on this task.
Additionally, our evaluation of LLMs is limited to those accessible via APIs. Models that can be run locally, such as those with 8B or 14B parameters, are not included. Preliminary experiments indicate that these smaller, locally runnable LLMs struggle to generate valid ShEx schemas and lack sufficient structured generation capabilities to reliably produce correct constraint syntax. Nevertheless, we believe that with further development, such models could become strong candidates for this task.
@@ Appendix B ShEx Specification
for tree= font=, grow=east, parent anchor=east, child anchor=west, anchor=west, align=center, s sep=5pt, , [Museum [wdt:P1174 [xsd:decimal [*] ] ] [wdt:P856 [IRI [*] ] ] [wdt:P17 [@<Country> [{1, 1}] ] ] [wdt:P31 [[ wd:Q33506 ] [{1, 1}] ] ] ]
@@ Appendix C Experimental Details
Listing 5 presents a sample input from the global setting experiments, derived from the few-shot examples, for the class ‘film award (Q4220917)’ and predicate ‘organizer (P664)’.
Table 6 details the performance of the gpt-4o-mini model across six different compositions of matching criteria, evaluating its precision, recall, and F1-score on both datasets. A clear trend observable for gpt-4o-mini is the consistent improvement in F1-scores as the matching criteria are relaxed. For instance, in the triples setting on the YAGO dataset, the F1-score climbs from 0.591 under ’Exact’ node constraint and ’Exact’ cardinality matching to 0.695 when ’Datatype’ abstraction is allowed for node constraints and cardinality is ’Loosened’. Similarly, on the more challenging Wikidata EntitySchema dataset, the global setting sees its F1-score rise from 0.312 (Exact/Exact) to 0.651 (Datatype/Loosened). This demonstrates that while gpt-4o-mini produces a solid number of perfectly accurate constraints, its output contains an even larger proportion of constraints that are valid under more flexible, practical interpretations, particularly when considering datatype abstractions.
Compared to DeepSeek-V3, gpt-4o-mini performs better under strict exact-matching conditions. However, as the matching criteria are relaxed, DeepSeek-V3 surpasses gpt-4o-mini. This indicates that while gpt-4o-mini excels at generating precisely accurate constraints, DeepSeek-V3 is more capable of producing constraints that fulfill the core functional requirements of ShEx, making it more practical for real-world applications.
| Settings | Matching Criteria | YAGO Schema | Wikidata EntitySchema | |||||
@@
| Matching Criteria | YAGO Schema | Wikidata EntitySchema | |||||
| Node Constraint | Cardinality | P | R | F1 | P | R | F1 |
| Exact | Exact | 0.111 | 0.081 | 0.092 | 0.106 | 0.099 | 0.096 |
| Subclass | Exact | 0.111 | 0.081 | 0.092 | 0.106 | 0.099 | 0.096 |
| Datatype | Exact | 0.536 | 0.393 | 0.445 | 0.211 | 0.189 | 0.188 |
| Exact | Loosened | 0.111 | 0.081 | 0.092 | 0.149 | 0.129 | 0.129 |
| Subclass | Loosened | 0.111 | 0.081 | 0.092 | 0.151 | 0.131 | 0.131 |
| Datatype | Loosened | 0.630 | 0.465 | 0.525 | 0.356 | 0.297 | 0.304 |

### v2

@@ 7 Conclusion
The size of the dataset is constrained by limited resources available for its construction. Creating high-quality shape schemas is a time-consuming process that requires multiple rounds of refinement by knowledge engineers. The limited size of the dataset makes it difficult to apply traditional ML approaches to the complete task.
Another limitation lies in the design of our evaluation metrics. While they capture the most essential features of schemas, they simplify schema structures and are therefore insufficient for handling more complex, deeply nested shapes. For example, our current formulation does not differentiate edit costs across specific constraint variations (e.g., predicates differing by the presence of the EXTRA keyword) and imported schemas.
Finally, our evaluation of LLMs is limited to those accessible via APIs. Models that can be run locally, such as those with 8B or 14B parameters, are not included. Preliminary experiments indicate that these locally runnable LLMs struggle to generate valid ShEx schemas and lack sufficient structured generation capabilities to reliably produce correct constraint syntax.
@@ Appendix C Experimental Details
Settings Matching Criteria YAGOS Node Constraint Cardinality P R F1 local Exact Exact 0.536 0.524 0.526 Subclass Exact 0.551 0.540 0.542 Datatype Exact 0.634 0.625 0.626 Exact Loosened 0.589 0.582 0.580 Subclass Loosened 0.604 0.598 0.596 Datatype Loosened 0.753 0.751 0.746 global Exact Exact 0.434 0.439 0.435 Subclass Exact 0.521 0.528 0.523 Datatype Exact 0.698 0.709 0.701 Exact Loosened 0.441 0.446 0.442 Subclass Loosened 0.532 0.541 0.535 Datatype Loosened 0.751 0.764 0.755 triples Exact Exact 0.535 0.505 0.510 Subclass Exact 0.554 0.523 0.528 Datatype Exact 0.630 0.591 0.600 Exact Loosened 0.581 0.550 0.554 Subclass Loosened 0.604 0.570 0.575 Datatype Loosened 0.759 0.720 0.725
Model YAGOS Acc (min) Acc (max) Acc sheXer 0.961 0.658 0.618 GPT-4o mini (local) 0.934 0.779 0.722 GPT-4o mini (global) 0.892 0.744 0.644 GPT-4o mini (triples) 0.942 0.684 0.658 DeepSeek-V3 (local) 0.880 0.786 0.729 DeepSeek-V3 (global) 0.969 0.768 0.737 DeepSeek-V3 (triples) 0.918 0.713 0.688 RDFShapeInduction (DT) 0.980 0.808 0.788 RDFShapeInduction (MLP) 0.827 0.689 0.579 RDFShapeInduction (RF) 0.989 0.809 0.798 RDFShapeInduction (GB) 0.976 0.797 0.773
Model YAGOS P R F1 GED NGED sheXer 0.085 0.056 0.066 43.67 0.616 GPT-4o mini (local) 0.669 0.631 0.649 14.89 0.218 GPT-4o mini (global) 0.342 0.294 0.314 27.56 0.398 GPT-4o mini (dt-global) 0.373 0.360 0.366 21.89 0.306 GPT-4o mini (gb-global) 0.356 0.353 0.354 22.33 0.315 GPT-4o mini (triples) 0.667 0.573 0.612 20.89 0.307 DeepSeek-V3 (local) 0.635 0.588 0.608 19.56 0.286 DeepSeek-V3 (global) 0.427 0.432 0.429 19.89 0.281 DeepSeek-V3 (dt-global) 0.466 0.448 0.456 19.22 0.269 DeepSeek-V3 (gb-global) 0.447 0.442 0.444 19.44 0.276 DeepSeek-V3 (triples) 0.603 0.488 0.535 24.11 0.348
@@
Examples of input prompts for the local, triples, and global settings are provided in Listings 5, 6, 7 for the class ‘film award (Q4220917)’.
Table 8 presents the results of DeepSeek-V3 on the YAGOS dataset across different matching criteria. Table 9 and 10 report the evaluation results for cardinality prediction and the performance of hybrid approaches on the YAGOS dataset. Table 11 and 12 show the performance of sheXer and GPT-4o mini models across six different compositions of matching criteria, evaluating their precision, recall, and F1-score on both datasets. A clear trend observable for GPT-4o mini is the consistent improvement in F1-scores as the matching criteria are relaxed. For instance, in the triples setting on the YAGO dataset, the F1-score climbs from 0.591 under ’Exact’ node constraint and ’Exact’ cardinality matching to 0.695 when ’Datatype’ abstraction is allowed for node constraints and cardinality is ’Loosened’. Similarly, on the more challenging Wikidata EntitySchema dataset, the global setting sees its F1-score rise from 0.312 (Exact/Exact) to 0.651 (Datatype/Loosened). This demonstrates that while GPT-4o mini produces a solid number of perfectly accurate constraints, its output contains an even larger proportion of constraints that are valid under more flexible, practical interpretations, particularly when considering datatype abstractions.
Compared to DeepSeek-V3, GPT-4o mini performs better under strict exact-matching conditions. However, as the matching criteria are relaxed, DeepSeek-V3 surpasses GPT-4o mini. This indicates that while GPT-4o mini excels at generating precisely accurate constraints, DeepSeek-V3 is more capable of producing constraints that fulfill the core functional requirements of ShEx, making it more practical for real-world applications.
| Matching Criteria | YAGO Schema | Wikidata EntitySchema | |||||
| Node Constraint | Cardinality | P | R | F1 | P | R | F1 |
| Exact | Exact | 0.111 | 0.081 | 0.092 | 0.106 | 0.099 | 0.096 |
| Subclass | Exact | 0.111 | 0.081 | 0.092 | 0.106 | 0.099 | 0.096 |
| Datatype | Exact | 0.536 | 0.393 | 0.445 | 0.211 | 0.189 | 0.188 |
| Exact | Loosened | 0.111 | 0.081 | 0.092 | 0.149 | 0.129 | 0.129 |
| Subclass | Loosened | 0.111 | 0.081 | 0.092 | 0.151 | 0.131 | 0.131 |
| Datatype | Loosened | 0.630 | 0.465 | 0.525 | 0.356 | 0.297 | 0.304 |
| Settings | Matching Criteria | YAGOS | WES | |||||